Two different fast-food restaurants charge different amounts for each burger meal. For Burger Queen, the cost is determined by the equation of 𝐶𝐶=12𝑏𝑏+30 where 𝐶𝐶 is the total cost of a burger meal in Rands and 𝑏𝑏is the amount of burger meals ordered. For MacDanny’s the total cost of a burger meal is calculated with the equation of 𝐶𝐶=30𝑏𝑏.

What is the cost of a burger meal in each case for a customer who ordered a burger meal?

4.1. Burger Queen? (2)

4.2. MacDanny? (2) What is the total cost of burger meals in each case for a customer who ordered

19 burger meals?

4.3. Burger Queen? (2)

4.4. MacDanny? (2)

4.5. Explain which equation you think is more realistic and why. (2)


1
Expert's answer
2021-07-16T15:28:50-0400

4.1.

A burger meal mean, b= 1

So for Burger Queen , "C=\frac{1}{2}*b+30=\frac{1}{2}*1+30=30.5"

4.2.

"C=30b =30*1=30"

4.3.

"C=\frac{1}{2}*b+30=\frac{1}{2}*19+30=39.5"

4.4.

"C=30b=30\u221719=570"

4.5.

C=30b It is more realistic

When you buy a burger costing 30.5 and 19 burgers costing 39.5 they are almost the same. It is not realistic.